OpenStudy (anonymous):

simply -2x cube root of 27 *3. My book says the answer is -6x cube root 3x. I can't figure out how to get the -6x?

OpenStudy (tkhunny):

\[-2\cdot \sqrt[3]{27} = -2\cdot 3\] It's not real clear what shoudl be done with that 'x'. It's just kind of floating about with no guidance.

OpenStudy (anonymous):

the full problem is -2x\[-2xsqrt[3]{81x}+4\sqrt[3]{3x ^{4}}\]

OpenStudy (anonymous):

I simplified the cube root of 81x to cube 27 * 3x. the cube of 27 is 3 so i don't understand how the book comes up with \[x6\sqrt[3]{3x}\]

OpenStudy (phi):

-2 out front

OpenStudy (anonymous):



OpenStudy (phi):

\[ -2x\sqrt[3]{81x}+4\sqrt[3]{3x ^{4}} \] \[ -2x \cdot 3 \sqrt[3]{3x} + 4 \cdot x\sqrt[3]{3x} \]

OpenStudy (phi):

that simplifies some more...

OpenStudy (anonymous):

ok, i understand, I was adding not multiplying. I knew it was something stupid and little just don't know why I couldn't figure that one out. Thanks so much.

OpenStudy (phi):

to finish this \[ -6x \sqrt[3]{3x}+4x\sqrt[3]{3x}= -2x\sqrt[3]{3x}\]